# Enhancing Trust in LLM-Generated Code Summaries with Calibrated Confidence Scores

### v2

@@ 3. Research Questions
In the first research question, we study whether an LLM’s confidence in its generated summary can be used to predict the quality of its generated summary. As previously established, we evaluate summary quality using metrics measuring BERTScore similarity to human-produced summaries. We first explore whether an LLM’s log-probabilities, can be directly used as a reliable confidence signal of the empirical likelihood of such similarity. For correctness, we transform the continuous BERTScore value into a binary indication of sufficient similarity, using thresholding: a summary is considered correct if it has a high enough BERTScore value. We try several approaches to obtaining a summary confidence measure from per-token LLM-generated probabilities, and several approaches to BERTScore thresholding. These are further discussed in Methods Section 4.4.
RQ 1. How well-calibrated are the LLM-produced confidence measures across all the tokens in the generated summary?
We do find that confidence measures directly calculated by averaging LLM per-token probabilities don’t work very well. We try rescaling, which is a method used to adjust a model’s confidences to better align with the observed frequencies of events. In this paper, we use Platt rescaling; please see discussion in Methods section.
@@ 4. Methodology
We begin with a description of the models & prompting strategies we used for the experiments; and then into our approach to calculating confidence, determining correctness, and calculating calibration.
@@ 4.1. Models
Code-Llama-70b. The Code Llama (Roziere et al., 2023), family of language models are tailored for coding tasks. It offers leading performance among open models. It has infilling capabilities, support for extensive input contexts, and zero-shot instruction-following abilities. The family includes foundation models, Python-specialized variants, and instruction-following models, each with 7B, 13B, 34B, and 70B parameters. Trained on 16k-token sequences, these models support inputs of up to 100k tokens. Notably, Code Llama achieved state-of-the-art results on various code benchmarks, at the time of experimentation. We use the larger 70B parameter model for our experiments, to get better performance.
DeepSeek-Coder-33b Instruct. The DeepSeek-Coder series introduces open-source code models ranging from 1.3B to 33B parameters, trained from scratch on a massive 2 trillion-token dataset (Guo et al., 2024). These models, pre-trained on high-quality project-level code corpora, utilize a fill-in-the-blank task with a 16K window to enhance code generation and infilling. Extensive evaluations demonstrate that DeepSeek-Coder achieves state-of-the-art performance among open-source code models across various benchmarks, outperforming closed-source models like Codex and GPT-3.5. We use the 33B paramater model for our study.
@@ 4.2. Prompting Methods for Code Summaries
Retrieval Augmented Few-shot Learning. Few-shot learning works well for both Natural Language Processing (Brown et al., 2020) and Software Engineering (Ahmed and Devanbu, 2022a; Nashid et al., 2023): we present the models with a few exemplars as query-answer pairs (method-comment pairs in this setup) and instruct the model to answer our final query (with just the method). Prior works show that few-shot learning performance can be improved using samples relevant to the query, identified using a retrieval algorithm (Ahmed et al., 2024; Nashid et al., 2023). Prior work also shows that the BM25 retrieval algorithm works better than vanilla few-shot learning for many SE tasks (e.g., code summarization, program repair, and assertion generation), so in this paper, we use few-shots retrieved with BM25.
Automatic Semantic Augmentation of Prompt (ASAP). Chain-of-thought (Wei et al., 2022) has been found to be effective for several tasks in NLP: providing the model with examples comprising explicit intermediate reasoning steps improves performance. For SE, Ahmed et al. (Ahmed et al., 2024), showed a workable approach to extracting such “intermediate steps” using static analysis algorithms; this approach enhances model performance. They used GitHub repository information, identifier type with scope, and dataflow information. Combined with BM25, this approach yielded state-of-the-art performance for code summarization tasks (Ahmed et al., 2024).
We therefore tried BM25 few-shot, also together with ASAP-based chain-of-thought, as our prompting strategy.
@@ 4.3. How to measure models’ confidence in a token sequence? (RQ1)
An LLM’s per-token probabilities reflect the likelihood of a token appearing in the same context in the (human-produced) training data. Therefore, we may expect a sequence-level probability to reflect the likelihood a generated summary is similar to a human-produced one. Still, the probabilities are calculated by the model per-token; so given a sequence of $N$ per-token probabilities, how should we compute a single representative probability for the entire sequence?
The combined sequence-level probability is (under an independence assumption) the product of per-token probabilities, and is usually very small, and varies with length. So, we normalize, using the $N$-th root of the product i.e. the geometric mean of probabilities. The arithmetic mean offers another option; however, since the product, not the sum, of individual probabilities better represents the probability of a sequence, we prefer the geometric mean. Empirically, we also consistently found better calibration when using the geometric mean over the arithmetic mean, since it less sensitive to large token-probability values (and thus reduces risk of overconfidence).
@@ 4.4. Evaluating similarity to human-produced summaries (RQ1)
In Section 2.3 we described the range of available metrics, and justified our focus on measures of the similarity to human-produced summaries. To evaluate the calibration of code summarization and answer each RQ, we need a binary notion of correctness which separates a good and bad summary. Our notion of correctness is sufficient similarity to human-produced summaries; this requires thresholding similarity metrics, which are real-valued quantities. We first determine these thresholds using available data from an existing human subject studies, and then next use these thresholds to study calibration performance of 3 models, on a much larger dataset used for studies of automated code summarization.
Dataset used for Setting Correctness Threshold We use the Haque et al. (Haque et al., 2022) dataset, which includes reference and generated summaries for 210 Java methods. For each Java method, it includes 3 human-assigned similarity ratings. Human raters select from 4 ratings ranging from ”Strongly Disagree” to ”Strongly Agree” indicating their agreement that the generated summary is similar to the reference. Haque et al. (Haque et al., 2022) used their data to analyze several different syntactic and semantic similarity metrics, and found that the semantic metric SentenceBERT offered the best correlation with human ratings.
Following their methodology, for each code method, we take the mean of the 3 human ratings to produce an aggregate, representative rating. We use this dataset to evaluate how accurately different automatic similarity metrics classify human agreement vs. disagreement on a summary’s similarity to the reference human-written summary. For our purposes, to qualify as a “correct” summary, the average human rating in the dataset must indicate at least an “Agree” rating.
Methodology for Selecting Thresholds We wanted thresholds for automated similarity metrics that are most consistent with this “Agree” level of human judgement of similarity. An optimal metric and threshold should classify all summaries that humans judge as similar as correct (perfect recall) while avoiding wrong classifications (perfect precision). As expected, all the metrics in Haque et al were imperfect; however, for illustration, with our criterion for correctness, we show an AUC-ROC curve for BERTScore (Figure 3) with an AUC of 0.86, which suggested that this metric, while not perfect, was a reasonable proxy for human judgement of similarity.
Given imperfect metrics, there are different ways to select different thresholds, based on the desired outcome. Lower thresholds would offer higher recall, but risk including summaries that may not resemble gold summaries. Higher thresholds would include fewer summaries, but ones more likely to resemble the gold ones. A cautious developer may want higher precision, lower recall; a less cautious developer might want some reasonable summary, even if lacking assurance of similarity to human summaries. For our purposes here, we would like to evaluate calibration in regards to different thresholds.
@@
Still, to obtain results for a range of settings, we wanted to obtain a workable set of thresholds over several different trade-offs, specifically: high recall, high precision, and optimal F1. To identify workable thresholds, we perform a grid search over the similarity metrics (from Haque et al), and threshold values, at 0.01 intervals to identify metric-threshold pairs meeting several criteria: 1) the highest recall while maintaining a high precision (¿ 0.9) 2) the highest precision with high recall (¿ 0.9) 3) Highest F1 score (harmonic mean of precision and recall). These metrics and thresholds respectively prioritize cautious, lenient, and balanced classifications of good vs. bad summaries, as viewed from the human judgement data. For each condition, the grid search respectively produces: 1) BERTScore at threshold 0.89 for high precision 2) Sentence-Bert at threshold 0.80 for high recall and 3) BERTScore at threshold 0.49 for optimal F1. We report results for all these settings. For illustration, Figure 3 (box) shows the two threshold values selected for BERTScore: for highest precision at at least 0.9 recall, the threshold is 0.89; for optimal F1, the threshold is 0.49.
@@ 4.5. Experimental Dataset
After setting the correctness thresholds for the BERTScore metric using the data set from (Haque et al., 2022), we needed a different dataset for the Calibration experiments. We use the Java and Python code summarization datasets from the CodeXGLUE (Lu et al., 2021) Benchmark for evaluating the calibration of the confidence values generated from LLM per-token probabilities. CodeXGLUE includes 14 datasets covering 10 Software Engineering tasks. Java and Python datasets contain 164,923 and 251,820 training samples, respectively. We use BM25 (Robertson et al., 2009) algorithms to retrieve relevant samples from these datasets for few-shotting. For the test set, we randomly select 5000 samples from each language. Only the training partition was used for BM25 retrieval. It should be noted that the CodeXGLUE has been de-duplicated, reducing the risk that an example from the test set would appear as a retrieved few-shot sample. It should also be noted that golden (human-written) summaries are available in this dataset; generated summaries can be compared with golden summaries to see if the semantic similarity score is above the given threshold.
@@ 4.6. Rescaling (RQ2)
Rescaling involves adjusting the predicted probabilities of a model to better align with the observed frequencies of events; this adjustment is typically done by fitting a low-parameter curve, using some of the data. Platt Scaling (Platt et al., 1999) is a commonly employed method in calibration, where a logistic regression model is fitted to the predicted probabilities alongside the corresponding true outcomes. Through this process, the model fine-tunes the predicted probabilities to more accurately reflect the true probabilities of events. While there are other rescaling approaches  (Guo et al., 2017; Zadrozny and Elkan, 2001, 2002), we use Platt scaling, since it has been successfully in prior software engineering work (Spiess et al., 2025).
To fit the Platt rescaling, we employ 5-fold cross-validation, where for each fold, we utilize the other 4 folds as training data and apply the learned regression to our target fold. We repeat this process for all other folds to obtain rescaled values for all samples. To accurately evaluate that rescaling is robust, we construct our folds such that each fold contains samples from different sets of repositories. This way, our test set is always pulled from different repositories than our training set; the reported Brier (and Skill) scores are calculated cumulatively over all folds for a more robust estimate.
@@ 4.7. Using the first tokens to compute confidences (RQ3)
Each token generated in a summary sequence successively constrains the next tokens in the sequence. For instance, the first few tokens might include an identifier, making the next verbs, prepositions, nouns etc most related to that identifier. Our initial study suggested that later tokens in a code summary have higher probabilities.
@@ 5. Results
We now present calibration results for summaries generated using few-shots retrieved by BM25. Table 1, Figure 4(a) and 4(b) show the raw calibration with respect to the optimal F1 thresholded BERTScore metric. Both the Brier score and ECE are undesirably high. Additionally, we observed a negative skill score for all cases. The Brier score ranges from 0.25 to 0.57, which are rather bad, indicating worse than random performance. For the CodeLlama-70b model, the model’s success rate is good for both Java (26%) and Python (23%). However, while these calibration values are poor, there is clearly some signal in the reliablity diagrams.
@@
We now study how calibration changes with rescaling. Table 2 and Figure 5 show significantly improved calibration for every configuration of model and language. All skill scores are positive, and both DeepSeek-Coder-33B and CodeLlama-70b exhibit very high skill scores (0.11) for Java and Python respectively. The Brier scores now range from 0.11 to 0.17 and the ECE ranges from 0.01 to 0.03, indicating good calibration. Although the skill score for GPT-3.5-Turbo is improved after rescaling, it is much lower than the other models. To summarize, the raw average token probability is not well-calibrated, but the rescaled average token probability is well-calibrated – with respect to thresholded BERTScore.
@@
As expected, when using a high precision correctness measure (BERTScore thresholded at 0.89), the model’s success rate is very low – virtually 0 for GPT-3.5-Turbo summaries of Python methods. For both Python and Java, every model except GPT-3.5-Turbo has excellent calibration with skill scores over 0.2. Since a lower success rate may produce a higher skill score (see 2), this high skill is partly due to the very low success rate; however, in addition to correctly producing low confidence for the many incorrect summaries, the Platt rescaler also learns to appropriately produce higher confidence for the few correct ones, thus resulting in higher skill.
When using a high recall correctness measure (SentenceBERT thresholded at 0.80), the success rate is significantly higher. We find similar calibration results to Table 3. Besides GPT-3.5-Turbo on Python methods, skill scores are over 0.08, indicating good calibration. Thus, even though more summaries are accepted as being above the threshold, the model’s output confidence (after rescaling) provides a usable indication as to whether the summary might adequately resemble a human-produced summary.
@@ 5.3. RQ3: Effect of Token Position on Confidence
We now present results on the relationship between token position and model confidence and its effect on calibration. Figure 6 shows that, as the position of a token in a code summary increases, the model’s confidence in the token also tends to increase. For Java summaries generated by Code-Llama-70b using retrieval-augmented few-shot prompting, after 20 tokens, the median token probability plateaus to over 0.9. Similar trends (where tokens late in the sequence exhibit high probabilities) are seen for every configuration of model, programming language, and prompting method. However, for some cases, this relationship is exhibited to a lesser extent. For example, in Python Summaries using GPT-3.5-Turbo, the median confidence computed with the first 5 tokens is already over 0.8. Although later tokens still tend to have even higher probabilities, since the mean probability over a sequence is already very large, their effect of excluding them isn’t so substantial.
Figure 7 shows the skill score achieved when only including the first $t$ tokens in the geometric mean. In figure 7(a), all first 7 tokens carry significant signal. However, including any more tokens reduces skill score since their probabilities are very large, smoothing the total mean towards high values which are not representative of uncertainty in the summary. In figure 7(b), only including early tokens in the geometric mean has little effect, primarily due to CodeLlama-70b generating short code summaries for Python. Since the proportion of generated code summaries with over 15 tokens is ¡10, only using the first e.g. 15 tokens has a diluted effect on skill score (which is measured across all summaries).
Figure 6. Distribution of token probabilities by its position in a generated summary. The y-axis shows the mean probability at a token position across all sequences. Summaries are of Java methods by CodeLlama-70b using Retrieval-Augmented Few-Shot prompting.
@@
Following Section 4.7, we find a statistically significant ($\alpha=0.01$) reduction in the Brier Scores (after Benjamini-Hochberg correction) for some but not all configurations of model, programming language, and prompting method. For Python summaries, we only find statistically significant reductions for CodeLlama-70b and DeepSeek-Coder-33b with few-shot prompting. For Java summaries, we find significant reductions for all models and prompting methods except DeepSeek-Coder-33b prompted with few-shot prompting.
Why is there a significant reduction in Brier Score in some cases and not others? Some configurations generate shorter summaries e.g. Figure 7(b). We find that code summaries generated for Python methods tend to be shorter than code summaries for Java. Since high probabilities expressed in later tokens aren’t as prevalent in shorter code summaries, they have a smaller effect on the mean confidence and thus calibration performance. Additionally, the selected hyperparameter $t$ might be poor. Particularly, for Java summaries generated with CodeLlama-70b and ASAP prompting, $t$=4 is chosen in some folds. By only using the first 4 tokens, significant information is lost. Finally, there might not be a substantial enough difference between probabilities early and late in the summary, such as for Python summaries generated by GPT-3.5-Turbo discussed at the start of 5.3.

### v3

@@ 3. Research Questions
In the first research question, we study whether an LLM’s confidence in its generated summary can be used to predict the quality of its generated summary. As previously established, we evaluate summary quality using metrics measuring BERTScore similarity to human-produced summaries. We first explore whether an LLM’s log-probabilities can be directly used as a reliable confidence signal of the empirical likelihood of such similarity. For correctness, we transform the continuous BERTScore value into a binary indication of sufficient similarity, using thresholding: a summary is considered correct if it has a high enough BERTScore value. We try several approaches to obtaining a summary confidence measure from per-token LLM-generated probabilities, and several approaches to BERTScore thresholding. These are further discussed in Methods Section 4.4.
RQ 1. Are the LLM-produced confidence measures over the tokens in the generated summary well-calibrated?
We find that confidence measures directly calculated by averaging LLM per-token probabilities are not well-calibrated. We try Platt rescaling, which is a method used to adjust a model’s confidences to better align with the observed frequencies of events.
@@ 4. Methodology
We describe the models & prompting strategies we used for the experiments; and then get into our approach to calculating confidence, determining correctness, and calculating calibration.
@@ 4.1. Models
Code-Llama-70b. The Code Llama (Roziere et al., 2023), family of language models are tailored for coding tasks. It offers leading performance among open models. It has infilling capabilities, support for extensive input contexts, and zero-shot instruction-following abilities. The family includes foundation models, Python-specialized variants, and instruction-following models, each with 7B, 13B, 34B, and 70B parameters. Trained on 16k-token sequences, these models support inputs of up to 100k tokens. Notably, Code Llama achieved state-of-the-art results on various code benchmarks at the time of experimentation. We use the larger 70B parameter model for our experiments.
DeepSeek-Coder-33b Instruct. The DeepSeek-Coder series introduces open-source code models ranging from 1.3B to 33B parameters, trained from scratch on a 2 trillion-token dataset (Guo et al., 2024). These models, pre-trained on high-quality project-level code corpora, utilize a fill-in-the-blank task with a 16K window to enhance code generation and infilling. Extensive evaluations demonstrate that DeepSeek-Coder achieves state-of-the-art performance among open-source code models across various benchmarks, outperforming closed-source models like Codex and GPT-3.5. Our study uses the 33B paramater model.
@@ 4.2. Prompting Methods for Code Summaries
Retrieval Augmented Few-shot Learning. Few-shot learning works well for both Natural Language Processing (Brown et al., 2020) and Software Engineering (Ahmed and Devanbu, 2022a; Nashid et al., 2023): we prompt models with 3 exemplars as query-answer pairs (method-comment pairs) and instruct the model to answer our final query (just the method). Prior works show that few-shot learning performance can be improved using exemplars relevant to the final query, identified using a retrieval algorithm (Ahmed et al., 2024; Nashid et al., 2023). Prior work also shows that the BM25 retrieval algorithm works better than vanilla few-shot learning for many SE tasks (e.g., code summarization, program repair, and assertion generation), so in this paper, we use few-shots retrieved with BM25.
Automatic Semantic Augmentation of Prompt (ASAP). Chain-of-thought (Wei et al., 2022) has been found to be effective for several tasks in NLP: providing the model with examples comprising explicit intermediate reasoning steps improves performance. For SE, (Ahmed et al., 2024) showed a workable approach (ASAP) to extracting such “intermediate steps” using static analysis algorithms; this approach enhances model performance. They used GitHub repository information, identifier type with scope, and dataflow information. Combined with BM25, this approach yielded state-of-the-art performance for code summarization tasks (Ahmed et al., 2024). We therefore also use BM25 few-shot together with ASAP-based chain-of-thought.
@@ 4.3. How to measure models’ confidence in a token sequence? (RQ1)
An LLM’s per-token probabilities reflect the likelihood of a token appearing in the same context in the (human-produced) training data. Therefore, we may expect a sequence-level probability to reflect the likelihood a generated summary is similar to a human-produced one. Still, the probabilities are calculated by the model per-token; so given a sequence of $N$ per-token probabilities, how should we compute a single representative probability?
The sequence-level probability (under an independence assumption) is the product of per-token probabilities. Since the product is usually very small and varies with sequence length, we normalize by taking its $N$-th root i.e. the geometric mean of probabilities. The arithmetic mean offers another option; however, since the product, not the sum, of individual probabilities better represents the probability of a sequence, we prefer the geometric mean. Empirically, we also consistently found better calibration when using the geometric mean over the arithmetic mean, since it less sensitive to large token-probability values (and thus reduces risk of overconfidence).
@@ 4.4. Evaluating similarity to human-produced summaries (RQ1)
In Section 2.3 we described the range of available metrics, and justified our focus on measures of the similarity to human-produced summaries. To evaluate the calibration of code summarization, we need a binary notion of correctness which separates a good and bad summary. Our notion of correctness is sufficient similarity to human-produced summaries; this requires thresholding similarity metrics, which are real-valued quantities. We first determine these thresholds using available data an existing human subject studies, and then use these thresholds to study calibration performance of 3 models, on a much larger dataset used for studies of automated code summarization.
Dataset used for Setting Correctness Threshold We use the  (Haque et al., 2022) dataset, which includes reference and generated summaries for 210 Java methods. For each Java method, it includes 3 human-assigned similarity ratings. Human raters select from 4 ratings ranging from ”Strongly Disagree” to ”Strongly Agree” indicating their agreement that the generated summary is similar to the reference.  (Haque et al., 2022) used their data to analyze several similarity metrics, and found that the semantic metric SentenceBERT offered the best correlation with human ratings.
Following their methodology, for each code method, we take the mean of the 3 human ratings to produce an aggregate rating. We use this dataset to evaluate how accurately different automatic similarity metrics classify human agreement vs. disagreement on a summary’s similarity to the reference summary. For our purposes, to qualify as a “correct” summary, the average human rating in the dataset must indicate at least an “Agree” rating.
Methodology for Selecting Thresholds We wanted thresholds for automated similarity metrics that are most consistent with this “Agree” level of human judgement of similarity.
@@
An optimal metric and threshold should classify all summaries that humans judge as similar as correct (perfect recall) while avoiding wrong classifications (perfect precision). As expected, all the metrics in Haque et al. were imperfect; however, for illustration, with our criterion for correctness, we show an AUC-ROC curve for BERTScore (Figure 3) with an AUC of 0.86, which suggested that this metric, while not perfect, was a reasonable proxy for human judgement of similarity.
Given imperfect metrics, there are different ways to select thresholds, based on the desired outcome. Lower thresholds would offer higher recall, but risk including summaries that may not resemble gold summaries. Higher thresholds would include fewer summaries, but ones more likely to resemble the gold ones. A cautious developer may want higher precision, lower recall; a less cautious developer might want some reasonable summary, even if lacking assurance of similarity to human summaries. For our purposes here, we would like to evaluate calibration for a set of thresholds over several different trade-offs, specifically: high recall, high precision, and optimal F1.
To identify workable thresholds, we perform a grid search over the similarity metrics (from Haque et al.), and threshold values, at 0.01 intervals to identify metric-threshold pairs meeting several criteria: 1) the highest recall while maintaining a high precision (¿ 0.9), 2) the highest precision with high recall (¿ 0.9), 3) the highest F1 score (harmonic mean of precision and recall). These metrics and thresholds respectively prioritize cautious, lenient, and balanced classifications of good vs. bad summaries, as viewed from the human similarity judgment data. For each condition, the grid search produces: 1) BERTScore at threshold 0.89 for high precision 2) Sentence-Bert at threshold 0.80 for high recall and 3) BERTScore at threshold 0.49 for optimal F1. We report results for all these settings. For illustration, Figure 3 shows the two threshold values selected for BERTScore.
@@ 4.5. Experimental Dataset
After setting the correctness thresholds for BERTScore using the dataset from (Haque et al., 2022), we needed a larger dataset for the calibration experiments. We use the Java and Python code summarization datasets from the CodeXGLUE Benchmark  (Lu et al., 2021) for evaluating the calibration of the confidence values generated from LLM per-token probabilities. CodeXGLUE includes 14 datasets covering 10 Software Engineering tasks. Java and Python datasets contain 164,923 and 251,820 training samples, respectively. We use BM25 (Robertson et al., 2009) algorithms to retrieve relevant samples from these datasets for few-shotting. For the test set, we randomly select 5000 samples from each language. Only the training partition was used for BM25 retrieval. It should be noted that the CodeXGLUE has been de-duplicated, reducing the risk that an example from the test set would appear as a retrieved few-shot sample. It should also be noted that reference (human-written) summaries are available in this dataset; generated summaries can be compared with reference summaries to see if the semantic similarity score is above the given threshold. While some human references in this dataset collected from open-source may be of poor quality (Fluri et al., 2007), the dataset as a whole captures content deemed as useful by actual developers. Among current evaluation approaches, the similarity to these docstrings captures this aspect best; casual human raters may lack the specific and deep knowledge the original developers have.
@@ 4.6. Rescaling (RQ2)
Rescaling involves adjusting the predicted probabilities of a model to better align with the observed frequencies of events. Platt Scaling (Platt et al., 1999) is a commonly employed method in calibration, where a logistic regression model is fitted to the predicted probabilities alongside the corresponding true outcomes. Through this process, the model transforms the predicted probabilities to more accurately reflect the true probabilities of events. While there are other rescaling approaches  (Guo et al., 2017; Zadrozny and Elkan, 2001, 2002), we use Platt scaling, since it has shown success in prior software engineering work (Spiess et al., 2025).
To fit the Platt rescaler, we employ 5-fold cross-validation, where for each fold, we utilize the other 4 folds as training data and apply the learned regression to our target fold. We repeat this process for all other folds to obtain rescaled values for all samples. To accurately evaluate that rescaling is robust, each fold only contains samples from different repositories. This way, test samples are always from different repositories than training samples; the reported Brier and Skill scores are calculated cumulatively over all folds.
@@ 4.7. Using the first tokens to compute confidences (RQ3)
Each token generated in a summary sequence successively constrains the next tokens in the sequence. For instance, the first few tokens might include an identifier, making the next verbs, prepositions, nouns, etc. most related to that identifier. Our initial study suggested that later tokens in a code summary have higher probabilities.
@@ 5. Result
We now present calibration results for summaries generated using few-shots retrieved by BM25. Table 1, Figure 4(a), and 4(b) show calibration results with respect to the optimal F1 thresholded BERTScore metric. Both the Brier score and ECE are undesirably high. Additionally, we observed a negative skill score for all cases. The Brier score ranges from 0.25 to 0.57, which indicate worse than random performance. While these calibration values are poor, the reliablity diagrams clearly show some signal.
We additionally note that model confidences show weak to moderate correlation to BERTScore. In the best case, Codellama-70b Python summaries produced a 0.45 Spearman correlation coefficient (Hollander et al., 2013). This result is expected since correlation is distinct from calibration. When viewing confidence as a probability indicating the likelihood of binary correctness, 90% confidence should not indicate 90% BERTScore; it should indicate the likelihood a generated summary is acceptable.
@@
We now report how calibration changes with rescaling. Table 2 and Figure 5 show significantly improved calibration for every configuration of model and programming language. All skill scores are positive, and both DeepSeek-Coder-33B and CodeLlama-70b exhibit very high skill scores (0.11) for Java and Python respectively. The Brier scores now range from 0.11 to 0.17 and the ECE ranges from 0.01 to 0.03, indicating good calibration. Although the skill score for GPT-3.5-Turbo is improved after rescaling, it is much lower than the other models. To summarize, the raw mean token probability is not well-calibrated, but the rescaled mean token probability is well-calibrated — with respect to thresholded BERTScore.
@@
As expected, when using a high precision correctness measure (BERTScore thresholded at 0.89), the model’s success rate is very low – virtually 0 for GPT-3.5-Turbo summaries of Python methods. For both Python and Java, every model except GPT-3.5-Turbo has excellent calibration with skill scores over 0.2. This high skill is partly due to the very low success rate (see equation 2). However, in addition to correctly producing low confidence for the many incorrect summaries, the Platt rescaler also learns to produce higher confidence for the few correct ones.
When using a high recall correctness measure (SentenceBERT thresholded at 0.80), the success rate is significantly higher. We find similar calibration results to Table 3. Besides GPT-3.5-Turbo on Python methods, skill scores are over 0.08, indicating good calibration. Thus, even though more summaries are accepted as being above the threshold, the model’s output confidence (after rescaling) provides a reliable indication as to whether the summary might adequately resemble a human-produced summary.
@@ 5.3. RQ3: Effect of Token Position on Confidence
We now present the relationship between token position and model confidence and its effect on calibration. Figure 6 shows that, as a token’s position in a Java code summary increases, CodeLlama-70b’s confidence in the token also tends to increase. After 20 tokens, the median token probability for each distribution plateaus to over 0.9. Similar trends (where tokens late in the sequence exhibit high probabilities) are seen for every configuration of model, programming language, and prompting method. However, in some cases, this relationship is exhibited to a lesser extent e.g. in GPT-3.5-Turbo-generated Python summaries, the median of token probabilities at positions 1-5 is already over 0.8. Although later tokens still tend to have even higher probabilities, excluding them during the mean probability calculation has a smaller effect since the mean probability of the cutoff sequence is already very large.
Figure 6. Distributions of token probabilities by their position in a generated summary. Summaries are of Java methods by CodeLlama-70b using Retrieval-Augmented Few-Shot prompting.
@@
Figure 7 shows the skill score achieved when only including the first $t$ tokens in the geometric mean. In Figure 7(a), all first 7 tokens carry significant signal. However, including any more tokens reduces skill score since their probabilities are very large, smoothing the total mean towards high values which are not representative of uncertainty in the summary. In Figure 7(b), only including early tokens in the geometric mean has little effect, primarily due to CodeLlama-70b generating short code summaries for Python (see the top plot in Figure 7(b)). Since the proportion of generated code summaries with over 15 tokens is only 9.41%, computing a mean using the first e.g. 15 tokens has a small effect on skill score (which is measured across all summaries).
Following Section 4.7, by using just the first t, instead of all tokens in computing the geometric mean, we find a statistically significant ($\alpha=0.01$) reduction in the Brier Scores (after Benjamini-Hochberg correction) for some but not all configurations of model, programming language, and prompting method. For Python summaries, we only find statistically significant reductions for CodeLlama-70b and DeepSeek-Coder-33b with few-shot prompting. For Java summaries, we find significant reductions for all models and prompting methods except DeepSeek-Coder-33b few-shot prompted.
Why is there a significant reduction in Brier Score in some cases and not others? Some configurations generate shorter summaries e.g. Figure 7(b). We find that code summaries generated for Python methods tend to be shorter than code summaries for Java. Additionally, the selected hyperparameter $t$ might be poor; for Java summaries generated with CodeLlama-70b and ASAP prompting, some folds use $t$=4 and may thus lose significant information. Finally, probabilities early in the summary may already be large e.g. GPT-3.5-Turbo-generated Python summaries discussed earlier in the section.